# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ III Generalization to Multi-Hop IRS-Assisted MIMO System
Tensor-based model formulation
In this section, we explore the $D$-IRS scheme connecting UTs and BSs, assuming unavailability of channels from UT to BS, from UT to the $i$-th IRS ($1<i\leq D$), from the $j$-th IRS ($1\leq j<D$) to BS, and between non-adjacent IRSs due to blockage or excessive path loss. In detail, we denote the UT-$1$-th IRS channel as $\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM}$, the $D$-th IRS-BS channel as $\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}}$, the channel from the $d$-th IRS to the $(d+1)$-th IRS as $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ for $d\in[D-1]$, and represent the $d$-th IRS as $\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}}$ for $d\in[D]$. Here, $\mathcal{S}_{d}(:,k,:)=\text{diag}(\bm{s}_{d}(k))\in\mathbb{C}^{N_{d}\times N_{d}}$, and the design of $\bm{s}_{d}(k)$ is same with (2). We define each element of the channel tensor $\mathcal{B}^{\star}_{D}\in\mathbb{C}^{LP\times K\times UM}$ as follows:
[…]
@@
where, in contrast to the multi-hop scheme [69], we assume the absence of a direct channel from the BS to the UT to streamline the analysis.
Note that each element $\mathcal{B}^{\star}_{D}(p,k,j)$ can still be conceptualized as an element of an order-$2D+1$ TT format tensor due to its multiplication form resembling (29). Specifically, to establish a connection between the channel tensor $\mathcal{B}^{\star}_{D}$ and the TT format tensor, we can reinterpret $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ as an order-$3$ tensor $\widetilde{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times 1\times N_{d}}$. Accordingly, we define ^$\widehat{\mathcal{B}}_{D}^{\star}$ as the result of a sequence of tensor contractions  [^4] Here, the tensor contraction operation $\mathcal{A}_{1}\times_{d}^{1}\mathcal{A}_{2}$ for $\mathcal{A}_{1}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ and $\mathcal{A}_{2}\in\mathbb{C}^{N_{1}\times M_{2}\times M_{3}}$ results in a new tensor $\mathcal{A}_{3}$ of size $N_{1}\times\cdots\times N_{d-1}\times N_{d+1}\times\cdots\times N_{D}\times M_{2}\times M_{3}$, with the $(n_{1},\dots,n_{d-1},n_{d+1},\dots,n_{D},m_{2},m_{3})$-th element being $\sum_{n_{d}}\mathcal{A}_{1}(n_{1},\dots,n_{d})\mathcal{A}_{2}(n_{d},m_{2},m_{3})$.: $\bm{B}_{D}\times_{2}^{1}\mathcal{S}_{D}\times_{3}^{1}\widetilde{\bm{B}}_{D-1}\times_{4}^{1}\cdots\times_{2D-1}^{1}\widetilde{\bm{B}}_{1}\times_{2D}^{1}\mathcal{S}_{1}\times_{2D+1}^{1}\bm{B}_{0}$. As a result, ^$\widehat{\mathcal{B}}_{D}^{\star}$ can be represented as an order-$2D+1$ TT format tensor with dimensions $LP\times K\times 1\times K\times\cdots\times 1\times K\times UM$ with TT ranks  [^5] The definition of TT ranks is provided in Appendix A. $(N_{D},N_{D},N_{D-1},N_{D-1},\dots,N_{1},N_{1})$. Further insights can be gained, revealing that
$\mathcal{B}^{\star}_{D}(p,k,j)$
={^$=\begin{cases}\widehat{\mathcal{B}}_{D}^{\star}(p,k_{1},1,k_{2},1,\dots,k_{D},j),&k_{1}=\dots=k_{D}=k,\\ 0,&\text{otherwise},\end{cases}$ (17)
where it becomes evident that $\mathcal{B}^{\star}_{D}$ emerges as a sampled outcome of ^$\widehat{\mathcal{B}}_{D}^{\star}$. To streamline the notation, we introduce the notation $\mathcal{B}^{\star}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}$. Consequently, we define the TT-based ToT regression model as following:
[…]
@@
^$\widehat{\mathcal{B}}_{D}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})-\widetilde{\mathcal{Y}}\|_{F}^{2},$ (19)
@@
[…]
@@
Error analysis
Before presenting the error analysis, we first ensure that the measurements satisfy the RIP condition. By following a similar derivation as provided in Appendix B, the RIP condition for $\mathcal{X}(\mathcal{B}_{D})$ can be directly obtained.
@@
Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable. Let $\delta_{UM}\in(0,1)$ be a positive constant. Then, for any channel tensor $\mathcal{B}^{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$, when the number of time slots satisfies
@@
with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
where $c$ and $C$ is a universal constant.
Similar to Theorem 1, Theorem 4 ensures that the number of time slots is solely contingent upon the total number of antennas (UM) in the UTs. This is attributed to the fact that the pilot signal matrix $\bm{X}$ is exclusively associated with the UT-$1$-th IRS channel $\bm{B}_{0}$. Then we can further analyze the recovery error ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$.
@@
(Upper bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}_{D}$ in (20) in which $\bm{B}_{d},d=0,\dots,D$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}$ for a positive constant $c_{1}$, we have
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (23)
where $N_{0}=UM$, $N_{D+1}=LP$ and  ^$\widehat{\mathcal{B}}_{D}$ is the solution to (19).
The proof has been provided in Appendix E. Theorem 5 ensures an optimal recovery error bound, given that the degree of freedom for the unknown variables in (19) is $O(LPN_{D}+UMN_{1}+\sum_{i=1}^{D-1}N_{d}N_{d+1})$. However, it is noteworthy that while increasing the number of IRSs can expand service coverage, the recovery error of channel estimation may also increase when the number of time slots $T$ is limited. Therefore, in practical implementation, a balance should be struck between service coverage and recovery error. Similar to the analysis of the minimax lower bound using the Gaussian ensemble design in Theorem 3, this result can also be extended to multi-hop IRS-assisted MIMO systems.
@@
(Minimax lower bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) In the context of the TT-based ToT regression model in (18), we assume a channel tensor $\mathcal{B}^{\star}_{D}$ from (20), where $\bm{B}_{d},d=0,\dots,D$ are full rank and $\min N_{d}\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
$\inf_{\widehat{\mathcal{B}}_{D}}\sup_{\mathcal{B}^{\star}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{\sum_{d=0}^{D}N_{d}N_{d+1}}{T}}\gamma\bigg{)}.$ (24)
Finally, akin to the analysis of a single-hop IRS-assisted MIMO system, it is imperative to address the following two special cases. $(i)$ First, when $\bm{B}_{d},d=0,\dots D$ are rank-$r_{d}$ matrices, by incorporating the covering number of low-rank matrices into (60) in Appendix E, we can derive
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
[…]
where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]\ \text{are unknown with }\text{rank}(\bm{B}_{d})=r_{d},d=0,\dots D,\ \text{and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$. Leveraging low-rank structures offers a potential avenue for improving error performance, as the recovery error may be reduced below (23) when the ranks $r_{d},d=0,\dots,D$ are small.
$(ii)$ Second, when $\mathcal{S}_{i},i\in[D]$, are unknown, introducing the covering number of $\{\mathcal{S}_{i}\}_{i=1}^{D}$ into the proof in Appendix E, similar to (23), we can arrive at
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
[…]
where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{N_{D+1}\times K\times N_{0}}:\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d=0,\dots,D,\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are unknown}\}$. In this analysis, we assume that all IRSs are equipped with active elements. However, in practice, when adjustments are required for only a subset of IRSs, denoted as $\mathcal{S}_{i}$ for $i\in\Omega$, where $\Omega\subseteq\{1,\dots,D\}$, the recovery error can be further reduced. Specifically, the term $\sum_{d=1}^{D}KN_{d}$ is replaced by $\sum_{i\in\Omega}KN_{i}$.
Discussion

### v2

@@ III Generalization to Multi-Hop IRS-Assisted MIMO System
III-A Tensor-based model formulation
In this section, we explore the $D$-IRS scheme connecting UTs and BSs, assuming unavailability of channels from UTs to BSs, from UTs to the $i$-th IRS ($1<i\leq D$), from the $j$-th IRS ($1\leq j<D$) to BSs, and between non-adjacent IRSs due to blockage or excessive path loss. In detail, we denote the UT-$1$-th IRS channel as $\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM}$, the $D$-th IRS-BS channel as $\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}}$, the channel from the $d$-th IRS to the $(d+1)$-th IRS as $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ for $d\in[D-1]$, and represent the $d$-th IRS as $\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}}$ for $d\in[D]$. Here, $\mathcal{S}_{d}(:,k,:)=\text{diag}(\bm{s}_{d}(k))\in\mathbb{C}^{N_{d}\times N_{d}}$, and the design of $\bm{s}_{d}(k)$ is same with (2). Here, we emphasize that, to simplify our theoretical analysis, the IRSs are arranged in a cascaded structure where only the first IRS interacts directly with the UEs, and only the last IRS connects to the BSs. All intermediate IRSs function purely as passive relays and do not directly interface with any BS or UE. A visual illustration can be found in [77, Figure 1], where, upon excluding the direct path, the remaining configuration accurately represents our model. We define each element of the channel tensor $\mathcal{B}^{\star}_{D}\in\mathbb{C}^{LP\times K\times UM}$ as follows:
$\mathcal{B}^{\star}_{D}(p,k,j)$ $\!\!\!\!=\!\!\!\!$ $\bm{B}_{D}(p,:)\Pi_{d=1}^{D-1}\mathcal{S}_{d+1}(:,k,:)\bm{B}_{d}$ (19)
@@ III-A Tensor-based model formulation
where we differ from the multi-hop scheme in [77] by assuming the absence of a direct BS-UT channel. In addition, we neglect both direct and reflected interference across distinct BS-UT links, effectively treating all paths as isolated. Furthermore, signal propagation and interference between different IRSs are neglected. These assumptions effectively decouple the communication paths, allowing us to concentrate on the highly nonconvex structure of the model in (19). To the best of our knowledge, a rigorous theoretical characterization of the recovery error for this model remains absent in the literature.
Note that each element $\mathcal{B}^{\star}_{D}(p,k,j)$ can still be conceptualized as an element of an order-$2D+1$ TT format tensor due to its multiplication form resembling (1) in the Appendix A. Specifically, to establish a connection between the channel tensor $\mathcal{B}^{\star}_{D}$ and the TT format tensor, we can reinterpret $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ as an order-$3$ tensor $\widetilde{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times 1\times N_{d}}$. Accordingly, we define $\widehat{\mathcal{B}}_{D}^{\star}$ as the result of a sequence of tensor contractions [^3] Here, the tensor contraction operation $\mathcal{A}_{1}\times_{d}^{1}\mathcal{A}_{2}$ for $\mathcal{A}_{1}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ and $\mathcal{A}_{2}\in\mathbb{C}^{N_{1}\times M_{2}\times M_{3}}$ results in a new tensor $\mathcal{A}_{3}$ of size $N_{1}\times\cdots\times N_{d-1}\times N_{d+1}\times\cdots\times N_{D}\times M_{2}\times M_{3}$, with the $(n_{1},\dots,n_{d-1},n_{d+1},\dots,n_{D},m_{2},m_{3})$-th element being $\sum_{n_{d}}\mathcal{A}_{1}(n_{1},\dots,n_{d})\mathcal{A}_{2}(n_{d},m_{2},m_{3})$.: $\bm{B}_{D}\times_{2}^{1}\mathcal{S}_{D}\times_{3}^{1}\widetilde{\bm{B}}_{D-1}\times_{4}^{1}\cdots\times_{2D-1}^{1}\widetilde{\bm{B}}_{1}\times_{2D}^{1}\mathcal{S}_{1}\times_{2D+1}^{1}\bm{B}_{0}$. As a result, $\widehat{\mathcal{B}}_{D}^{\star}$ can be represented as an order-$2D+1$ TT format tensor with dimensions $LP\times K\times 1\times K\times\cdots\times 1\times K\times UM$ with TT ranks [^4] The definition of TT ranks is provided in the Appendix A. $(N_{D},N_{D},N_{D-1},N_{D-1},\dots,N_{1},N_{1})$. Further insights can be gained, revealing that
$\hskip-22.76228pt\mathcal{B}^{\star}_{D}(p,k,j)$
$\hskip-22.76228pt=\begin{cases}\widehat{\mathcal{B}}_{D}^{\star}(p,k_{1},1,k_{2},1,\dots,k_{D},j),&k_{1}=\dots=k_{D}=k,\\ 0,&\text{otherwise},\end{cases}$ (21)
where it becomes evident that $\mathcal{B}^{\star}_{D}$ emerges as a sampled outcome of $\widehat{\mathcal{B}}_{D}^{\star}$. To streamline the notation, we introduce the notation $\mathcal{B}^{\star}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}$. Consequently, we define the TT-based ToT regression model as following:
$\widetilde{\mathcal{Y}}$ $\!\!\!\!=\!\!\!\!$ $\mathcal{X}(\mathcal{B}^{\star}_{D})+\mathcal{W}=\mathcal{B}^{\star}_{D}\times_{3}^{1}\bm{X}+\mathcal{W}$ (22)
$\!\!\!\!=\!\!\!\!$ $[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\times_{3}^{1}\bm{X}+\mathcal{W}.$
Here, we emphasize that the TT-based ToT regression model is equivalent to (20), and serves to simplify the subsequent error analysis.
@@
$\widehat{\mathcal{B}}_{D}=\operatorname*{\text{arg~min}}_{\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})-\widetilde{\mathcal{Y}}\|_{F}^{2},$ (23)
@@
$\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$ $\!\!\!\!\!\!=\!\!\!\!\!\!$ $\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:$ (24)
@@
III-B Optimal error analysis
Before presenting the error analysis, we first ensure that the measurements satisfy the RIP condition. By following a similar derivation as provided in Appendix C, the RIP condition for $\mathcal{X}(\mathcal{B}_{D})$ can be directly obtained.
@@ III-B Optimal error analysis
A linear operator $\mathcal{X}:\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ is said to satisfy the $UM$-RIP with a constant $\delta_{UM}$ if
holds for any channel tensor $\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$. We now consider two cases:
  • Suppose each element of the pilot signal matrix $\bm{X}$ is an i.i.d. complex-valued subgaussian random variable. For a positive constant $\delta_{UM}\in(0,1)$, when the number of time slots satisfies
@@
    with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP with positive constants $c$ and $C$.
  • Suppose $\frac{1}{\sqrt{T}}\bm{X}$ is the normalized DFT matrix, when
    $T\geq UM$ (27)
    is satisfied, we have $\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})\|_{F}^{2}\!=\!\|\mathcal{B}_{D}\|_{F}^{2}$ with $\delta_{UM}=0$.
Similar to Theorem 1, Theorem 4 ensures that the number of time slots is solely contingent upon the total number of antennas (UM) in the UTs. This is attributed to the fact that the pilot signal matrix $\bm{X}$ is exclusively associated with the UT-$1$-th IRS channel $\bm{B}_{0}$. Then we can further analyze the recovery error $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$.
@@
(Upper bound of $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}_{D}$ in (24) in which $\bm{B}_{d},d=0,\dots,D$ are full rank, when $\mathcal{X}$ satisfies the $UM$-RIP and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}$ for a positive constant $c_{1}$, the solution in (23) satisfies
$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (28)
where $N_{0}=UM$, $N_{D+1}=LP$ and  $\widehat{\mathcal{B}}_{D}$ is the solution to (23).
The proof has been provided in Appendix G. Theorem 5 provides an upper bound on the recovery error derived from a constrained least squares minimization, which scales with the total degrees of freedom of the unknown variables, i.e., $LPN_{D}+UMN_{1}+\sum_{i=1}^{D-1}N_{d}N_{d+1}$. However, it is noteworthy that while increasing the number of IRSs can expand service coverage, the recovery error of channel estimation may also increase when the number of time slots $T$ is limited. Therefore, in practical implementation, a balance should be struck between service coverage and recovery error. A discussion of special cases analogous to the single-hop IRS-assisted MIMO system is presented in Appendix H.
Building on the upper bound derived from constrained least squares minimization, we now establish a minimax lower bound to assess its optimality.
@@
(Minimax lower bound of $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) In the context of the TT-based ToT regression model in (22), we assume a channel tensor $\mathcal{B}^{\star}_{D}$ from (24), where $\bm{B}_{d},d=0,\dots,D$ are full rank and $\min N_{d}\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assume that the noise $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, and that the pilot matrix $\bm{X}$ is chosen as either (i) an i.i.d. $\mathcal{C}\mathcal{N}(0,1)$ matrix or (ii) a normalized DFT matrix $\frac{1}{\sqrt{T}}\bm{X}$. Under these two choices, we obtain
$\inf_{\widehat{\mathcal{B}}_{D}}\sup_{\mathcal{B}^{\star}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}=\Omega\bigg(\sqrt{\frac{\sum_{d=0}^{D}N_{d}N_{d+1}}{T}}\gamma\bigg).$ (29)
Our minimax lower bound for the TT-based ToT regression model in (22) demonstrates that the recovery error upper bound is tight and optimal up to the log term.
Figure 2: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system (a) for different NN with $UM=LP=K=10$, $T=20$ and $\gamma^{2}=10^{-6}$, (b) for different $T$ with $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (c) for different $\gamma^{2}$ with $UM=LP=K=N=10$ and $T=20$.
III-C Discussion